OpenStudy (anonymous):

I have the answer can somone please check me .... Find (dy/dx) x^3+y^3=3x

OpenStudy (nathan917):

Hi When differentiating a function of y with respect to x, take the derivative with respect to y and multiply it by y'. Also, remember to use the product rule. x^3 + y^3 = 6xy d/dx[x^3 + y^3] = d/dx[6xy] 3x^2 + (3y^2)y' = 6[x' * y + y' * x] 3x^2 + (3y^2)y' = 6y + 6xy' Solve for y' now. 3x^2 + (3y^2)y' = 6y + 6xy' 3x^2 + (3y^2)y' - 6xy' = 6y (3y^2)y' - 6xy' = 6y - 3x^2 (y')(3x^2 - 6) = 6y - 3x^2 y' = (6y - 3x^2)/(3x^2 - 6) y' = (2y - x^2)/(x^2 - 2) http://ca.answers.yahoo.com/question/index?qid=20101008020900AAMdNZM

OpenStudy (anonymous):

so then what would the answer be I;m confused cause that is not what I got.